# TikTok’s recommendations skewed towards Republican content during the 2024 U.S. presidential race

### v1

@@ 10.1. None of the above topics
Using the above prompt, the topics discussed in the 7,767 ideologically stanced TikTok videos in our dataset were identified. Supplementary Table S10 below describes the number of videos falling under each category, split by the video’s ideological stance.
@@
Table S10: The number of videos of each political classification on a certain topic.
@@
To verify that the Republican skew observed in our experiment are not due to differences in the engagement metrics of Republican and Democrat videos or channels, we build a series of counterfactual-models to predict expected video recommendations based on these engagement metrics. Specifically, we aim to compute differences between observed recommendation rates and predicted recommendation rates when sampling from the dataset of videos based on a series of different metrics. The metrics of interest include several video-level metrics, such as the number of plays/views a video receives, as well as its number of shares, comments, and likes. In addition to video-level metrics, we also collect channel level metrics such as a channel’s cumulative number of likes, followers, videos, as well as whether the channel is verified or not. Given that channel verification status is a binary attribute (True or False), we apply a value of 0.1 to channels who are not verified, and 0.9 to channels who are verified. For robustness, we also test all values between 0.5 and 1.0 in increments of 0.05 for verified channels (and inversely, 0.5 to 0.0 for unverified channels), and find similar results (see Supplementary Table S18).
@@
Lastly, for each of the aforementioned metrics, we repeat the sampling processes while taking into account the potential favorability of recent content. Specifically, we inversely scale each metric in question with the time since its publishing date, once linearly, and once exponentially. In total, this amounts to 39 different counterfactual models with which we compare observed recommendation rates and predicted recommendation rates. Supplementary Table S14 details the ideological skew computed by each model, as well as t-test results between the observed skew and expected skew based on each model. Supplementary Tables S15 and  S16 detail these results for models using only positive-partisan or negative-partisan videos, respectively.
Additional Supplementary Tables
@@
Table S11: The number of successful experimental runs per condition per week.
@@
Table S12: For each experimental condition, the proportion of videos that are political, and of those, the proportion of videos that pertain to the US elections or major political candidates. For each set of videos, the distribution of their ideological stances.
@@
Table S13: The mean ideological content seen by bots of different conditioning in each state over time.
@@
Table S14: Independent t-test results comparing ideological content of bots of different conditioning in each state over time. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
Video Comment Count (Linear) 0.062 -12.242***
Video Like Count (Linear) 0.061 -12.483***
Video Play Count (Linear) 0.065 -11.816***
Video Share Count (Linear) 0.075 -11.471***
Channel Verified (Linear) 0.042 -12.757***
Channel Follower Count (Linear) 0.051 -11.752***
Channel Like Count (Linear) 0.028 -11.254***
Channel Video Count (Linear) 0.033 -12.457***
Video Likes per Recomm. (Linear) 0.056 -12.192***
Video Comments per Recomm. (Linear) 0.059 -12.076***
Video Combined (Linear) 0.064 -12.184***
Channel Combined (Linear) 0.042 -12.82***
Full Combined (Linear) 0.044 -12.591***
Video Comment Count (Exponential) 0.058 -11.569***
Video Like Count (Exponential) 0.057 -11.903***
Video Play Count (Exponential) 0.060 -11.242***
Video Share Count (Exponential) 0.066 -11.026***
Channel Verified (Exponential) 0.049 -12.317***
Channel Follower Count (Exponential) 0.052 -12.141***
Channel Like Count (Exponential) 0.039 -13.551***
Channel Video Count (Exponential) 0.041 -13.583***
Video Likes per Recomm. (Exponential) 0.051 -12.277***
Video Comments per Recomm. (Exponential) 0.052 -12.111***
Video Combined (Exponential) 0.059 -11.427***
Channel Combined (Exponential) 0.047 -12.637***
Full Combined (Exponential) 0.048 -12.479***
Video Comment Count (No scaling) 0.083 -17.44***
Video Like Count (No scaling) 0.079 -23.173***
Video Play Count (No scaling) 0.096 -18.56***
Video Share Count (No scaling) 0.098 -17.022***
Channel Verified (No scaling) 0.058 -16.854***
Channel Follower Count (No scaling) 0.075 -17.368***
Channel Like Count (No scaling) 0.032 -21.316***
Channel Video Count (No scaling) 0.042 -20.746***
Video Likes per Recomm. (No scaling) 0.071 -19.435***
Video Comments per Recomm. (No scaling) 0.077 -16.223***
Video Combined (No scaling) 0.088 -18.848***
Chan l Combined (No scaling) 0.053 -18.018***
Full Combined (No scaling) 0.057 -18.114***
@@
Table S15: The ideological skew observed in the experiments, as well as the ideological skew computed through counterfactual models which sample videos based on a given attribute. The right-most column denotes independent t-test results between the expected ideological skew and that observed by the bots. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
Video Comment Count (Linear) -0.0068 -4.421***
Video Like Count (Linear) -0.0059 -4.125***
Video Play Count (Linear) -0.0037 -3.656***
Video Share Count (Linear) -0.0031 -3.757***
Channel Verified (Linear) -0.0180 -6.181***
Channel Follower Count (Linear) -0.0094 -4.765***
Channel Like Count (Linear) -0.0152 -5.379***
Channel Video Count (Linear) -0.0118 -4.915***
Video Likes per Recomm. (Linear) -0.0135 -5.332***
Video Comments per Recomm. (Linear) -0.01247 -5.093***
Video Combined (Linear) -0.0055 -4.087***
Channel Combined (Linear) -0.0156 -5.779***
Full Combined (Linear) -0.0147 -5.494***
Video Comment Count (Exponential) -0.0074 -4.126***
Video Like Count (Exponential) -0.0069 -4.01***
Video Play Count (Exponential) -0.0043 -3.445***
Video Share Count (Exponential) -0.0047 -3.57***
Channel Verified (Exponential) -0.0150 -5.693***
Channel Follower Count (Exponential) -0.0108 -4.765***
Channel Like Count (Exponential) -0.0126 -5.321***
Channel Video Count (Exponential) -0.0089 -4.44***
Video Likes per Recomm. (Exponential) -0.0123 -5.019***
Video Comments per Recomm. (Exponential) -0.0118 -5.045***
Video Combined (Exponential) -0.0060 -3.816***
Channel Combined (Exponential) -0.0138 -5.47***
Full Combined (Exponential) -0.0128 -5.217***
Video Comment Count (No scaling) 0.0094 -1.878
Video Like Count (No scaling) 0.0076 -2.537*
Video Play Count (No scaling) 0.0153 -0.19
Video Share Count (No scaling) 0.0114 -1.51
Channel Verified (No scaling) -0.0136 -6.531***
Channel Follower Count (No scaling) 0.0043 -3.225**
Channel Like Count (No scaling) -0.0026 -4.782***
Channel Video Count (No scaling) 0.0066 -2.191*
Video Likes per Recomm. (No scaling) -0.0039 -5.501***
Video Comments per Recomm. (No scaling) 0.0009 -4.195***
Video Combined (No scaling) 0.0107 -1.608
Channel Combined (No scaling) -0.0095 -5.85***
Full Combined (No scaling) -0.0060 -5.209***
@@
Table S16: The ideological skew observed in the experiments, as well as the ideological skew computed through counterfactual models which sample videos based on a given attribute when only considering positive-partisanship videos (Pro Democrat or Pro Republican). The right-most column denotes independent t-test results between the expected ideological skew and that observed by the bots. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
Video Comment Count (Linear) 0.0701 -11.868***
Video Like Count (Linear) 0.0675 -12.952***
Video Play Count (Linear) 0.0693 -12.49***
Video Share Count (Linear) 0.0792 -11.332***
Channel Verified (Linear) 0.0598 -12.692***
Channel Follower Count (Linear) 0.0616 -11.402***
Channel Like Count (Linear) 0.0438 -11.521***
Channel Video Count (Linear) 0.0447 -12.743***
Video Likes per Recomm. (Linear) 0.0704 -12.202***
Video Comments per Recomm. (Linear) 0.0723 -11.488***
Video Combined (Linear) 0.0704 -12.206***
Channel Combined (Linear) 0.0580 -12.397***
Full Combined (Linear) 0.0584 -12.49***
Video Comment Count (Exponential) 0.0659 -12.483***
Video Like Count (Exponential) 0.0640 -13.136***
Video Play Count (Exponential) 0.0659 -12.539***
Video Share Count (Exponential) 0.0706 -12.169***
Channel Verified (Exponential) 0.0649 -12.354***
Channel Follower Count (Exponential) 0.0622 -12.933***
Channel Like Count (Exponential) 0.0525 -13.805***
Channel Video Count (Exponential) 0.0499 -14.187***
Video Likes per Recomm. (Exponential) 0.0635 -12.753***
Video Comments per Recomm. (Exponential) 0.0652 -12.445***
Video Combined (Exponential) 0.0657 -12.596***
Channel Combined (Exponential) 0.0620 -12.705***
Full Combined (Exponential) 0.0623 -12.67***
Video Comment Count (No scaling) 0.0756 -16.235***
Video Like Count (No scaling) 0.0726 -24.122***
Video Play Count (No scaling) 0.0812 -20.731***
Video Share Count (No scaling) 0.0878 -15.986***
Channel Verified (No scaling) 0.0711 -17.35***
Channel Follower Count (No scaling) 0.0704 -18.685***
Channel Like Count (No scaling) 0.0359 -24.195***
Channel Video Count (No scaling) 0.0353 -24.26***
Video Likes per Recomm. (No scaling) 0.0751 -18.194***
Video Comments per Recomm. (No scaling) 0.0749 -15.136***
Video Combined (No scaling) 0.0775 -18.4***
Channel Combined (No scaling) 0.0620 -18.986***
Full Combined (No scaling) 0.0630 -19.682***
@@
Table S17: The ideological skew observed in the experiments, as well as the ideological skew computed through counterfactual models which sample videos based on a given attribute when only considering negative-partisanship videos (Anti Democrat or Anti Republican). The right-most column denotes independent t-test results between the expected ideological skew and that observed by the bots. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
Weight = 0.5 (Linear) 0.0606 -14.455***
Weight = 0.55 (Linear) 0.0585 -14.273***
Weight = 0.6 (Linear) 0.0575 -14.304***
Weight = 0.65 (Linear) 0.0549 -14.245***
Weight = 0.7 (Linear) 0.0541 -14.376***
Weight = 0.75 (Linear) 0.0513 -14.292***
Weight = 0.8 (Linear) 0.0506 -14.25***
Weight = 0.85 (Linear) 0.0495 -13.895***
Weight = 0.9 (Linear) 0.0465 -14.266***
Weight = 0.95 (Linear) 0.0450 -14.213***
Weight = 1.0 (Linear) 0.0263 -14.453***
Weight = 0.5 (Exponential) 0.0624 -12.716***
Weight = 0.55 (Exponential) 0.0606 -12.91***
Weight = 0.6 (Exponential) 0.0596 -13.047***
Weight = 0.65 (Exponential) 0.0577 -13.068***
Weight = 0.7 (Exponential) 0.0579 -13.174***
Weight = 0.75 (Exponential) 0.0557 -13.381***
Weight = 0.8 (Exponential) 0.0541 -13.417***
Weight = 0.85 (Exponential) 0.0522 -13.526***
Weight = 0.9 (Exponential) 0.0497 -13.805***
Weight = 0.95 (Exponential) 0.0479 -13.881***
Weight = 1.0 (Exponential) 0.0306 -14.022***
Weight = 0.5 (No scaling) 0.0864 -16.33***
Weight = 0.55 (No scaling) 0.0838 -16.633***
Weight = 0.6 (No scaling) 0.0808 -16.708***
Weight = 0.65 (No scaling) 0.0768 -16.652***
Weight = 0.7 (No scaling) 0.0757 -16.297***
Weight = 0.75 (No scaling) 0.0719 -16.855***
Weight = 0.8 (No scaling) 0.0678 -16.406***
Weight = 0.85 (No scaling) 0.0648 -17.296***
Weight = 0.9 (No scaling) 0.0613 -16.645***
Weight = 0.95 (No scaling) 0.0573 -17.438***
Weight = 1.0 (No scaling) 0.0502 -16.852***
@@
Table S18: The ideological skew observed in the experiments, as well as the ideological skew computed through counterfactual models which sample videos based on channel verification status with varying weights given to a channels verification status. The right-most column denotes independent t-test results between the expected ideological skew and that observed by the bots. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
Table S19: The 20 most common hashtags for Democrat and Republican conditioning videos.
@@
Table S20: The proportion of Democrat and Republican videos on a given topic, the difference in these proportions, and chi-squared tests comparing these proportions. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
Table S21: VIF values for control variables in logistic regression considering all videos viewed during experiment.
@@
Table S22: VIF values for control variables in logistic regression considering videos by top Democrat and Republican channels viewed during experiment.
@@
Week of experiment -0.0494 0.9518 0.9431 0.9606 p <0.001

### v2

@@ 10.1. None of the above topics
Using the above prompt, the topics discussed in the 7,767 ideologically stanced TikTok videos in our dataset were identified. Table 13 below describes the number of videos falling under each category, split by the video’s ideological stance.
@@
Table 13: The number of videos of each political classification on a certain topic.
@@
To verify that the Republican skew observed in our experiment are not due to differences in the engagement metrics of Republican and Democrat videos or channels, we build a series of counterfactual-models to predict expected video recommendations based on these engagement metrics. Specifically, we aim to compute differences between observed recommendation rates and predicted recommendation rates when sampling from the dataset of videos based on a series of different metrics. The metrics of interest include several video-level metrics, such as the number of plays/views a video receives, as well as its number of shares, comments, and likes. In addition to video-level metrics, we also collect channel level metrics such as a channel’s cumulative number of likes, followers, videos, as well as whether the channel is verified or not. Given that channel verification status is a binary attribute (True or False), we apply a value of 0.1 to channels who are not verified, and 0.9 to channels who are verified. For robustness, we also test all values between 0.5 and 1.0 in increments of 0.05 for verified channels (and inversely, 0.5 to 0.0 for unverified channels), and find similar results (see Supplementary Table 8).
@@
Lastly, for each of the aforementioned metrics, we repeat the sampling processes while taking into account the potential favorability of recent content. Specifically, we inversely scale each metric in question with the time since its publishing date, once linearly, and once exponentially. In total, this amounts to 39 different counterfactual models with which we compare observed recommendation rates and predicted recommendation rates. Supplementary Table 5 details the ideological skew computed by each model, as well as t-test results between the observed skew and expected skew based on each model. Supplementary Tables 6 and  7 detail these results for models using only positive-partisan or negative-partisan videos, respectively.
Survey
We administered a pre-registered survey with a sample of 1,008 U.S.-based TikTok users to assess whether individuals had noticed changes to the content of their TikTok feeds, particularly political content, over the past year. The survey was preregistered on OSF (https://osf.io/udywb/) and was deemed exempt by the authors’ institutional review board (IRB Protocol Number: HRPP-2025-69).
Participants were recruited via the online platform Prolific and screened to ensure they resided in the United States and were active users of TikTok. The survey consisted of two parts: (1) a series of open-ended text entry questions, and (2) a series of structured, scale-based questions. Open-ended items asked participants whether they had noticed any changes to the content on their TikTok feed in general, any changes to political content specifically, and whether the tone of political content had become more positive or negative. Responses to these questions were manually coded by the first author to determine whether participants explicitly referenced changes to political content and, if so, whether they described seeing more Republican-aligned or Democratic-aligned content.
Structured questions asked participants to rate, on a 0–10 scale, the extent to which their feed had shifted toward Democratic or Republican content, become more positive or negative in tone, or featured more political content they agreed or disagreed with.
For each survey item, we conducted separate linear regression analyses that included participant political affiliation and demographic covariates (age, gender, race, and education level) as predictors. Full regression results are presented in Supplementary Tables 20 and 21 for the open-ended and structured questions, respectively. Participant demographic characteristics are summarized in Supplementary Table 22, and the complete survey instrument is available in the Supplementary Materials under the Survey Materials section.
Author Contributions
Acknowledgments
We gratefully acknowledge Jason Greenfield, Kevin Munger, Manoel Horta Ribeiro, Chris Schwartz, Josh Tucker, and the Center for Social Media and Politics’ Data Science Reading Group for invaluable feedback during the preparation and revision of this manuscript.
Competing Interests
The authors declare no competing interests.
Data and Code Availability
Additional Information
Supplementary Information is available for this paper. Correspondence and requests for materials should be addressed to Yasir Zaki.
    Supplementary Tables
@@ Acknowledgments
    Supplementary Table 1: The number of successful experimental runs per condition per week.
@@
    Supplementary Table 2: For each experimental condition, the proportion of videos that are political, and of those, the proportion of videos that pertain to the US elections or major political candidates. For each set of videos, the distribution of their ideological stances.
@@
    Supplementary Table 3: The mean ideological content seen by bots of different conditioning in each state over time.
@@
    Supplementary Table 4: Independent t-test results comparing ideological content of bots of different conditioning in each state over time. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
    Comment Count (Exponential) -0.024 -11.922***
    Comment Count (Linear) -0.037 -12.479***
    Comment Count (No scaling) -0.009 -11.191***
    Comments per recommendation (Exponential) -0.05 -12.972***
    Comments per recommendation (Linear) -0.059 -13.373***
    Comments per recommendation (No scaling) -0.04 -12.403***
    Combined Author Score (Exponential) -0.059 -13.343***
    Combined Author Score (Linear) -0.065 -13.607***
    Combined Author Score (No scaling) -0.052 -13.067***
    Combined Likes, Shares, Plays Score (Exponential) 0.008 -10.162***
    Combined Likes, Shares, Plays Score (Linear) -0.005 -10.884***
    Combined Likes, Shares, Plays Score (No scaling) 0.032 -8.949***
    Combined Video Score (Exponential) -0.018 -11.63***
    Combined Video Score (Linear) -0.035 -12.36***
    Combined Video Score (No scaling) -0.001 -10.801***
    Like Count (Exponential) 0.002 -10.438***
    Like Count (Linear) -0.01 -11.126***
    Like Count (No scaling) 0.027 -9.238***
    Video Length (Exponential) -0.068 -13.926***
    Video Length (Linear) -0.08 -14.356***
    Video Length (No scaling) -0.07 -14.111***
    (Likes + Comments + Shares) / Plays (Exponential) -0.029 -12.142***
    (Likes + Comments + Shares) / Plays (Linear) -0.043 -12.788***
    (Likes + Comments + Shares) / Plays (No scaling) -0.009 -11.263***
    Channel followers (Exponential) -0.031 -12.282***
    Channel followers (Linear) -0.042 -12.672***
    Channel followers (No scaling) -0.015 -11.479***
    Full Combined score (Exponential) -0.051 -13.12***
    Full Combined score (Linear) -0.061 -13.55***
    Full Combined score (No scaling) -0.044 -12.815***
    Channel cumul. likes (Exponential) -0.05 -13.15***
    Channel cumul. likes (Linear) -0.057 -13.321***
    Channel cumul. likes (No scaling) -0.042 -12.925***
    Likes per recommendation (Exponential) -0.028 -12.114***
    Likes per recommendation (Linear) -0.042 -12.796***
    Likes per recommendation (No scaling) -0.009 -11.243***
    Video plays (Exponential) 0.01 -9.976***
    Video plays (Linear) -0.004 -10.773***
    Video plays (No scaling) 0.036 -8.714***
    Video shares (Exponential) 0.006 -10.341***
    Video shares (Linear) -0.005 -10.972***
    Video shares (No scaling) 0.029 -9.242***
    Channel video count (Exponential) -0.042 -12.728***
    Channel video count (Linear) -0.052 -13.077***
    Channel video count (No scaling) -0.03 -12.294***
    Channel verified (Exponential) -0.059 -11.519***
    Channel verified (Linear) -0.065 -11.803***
    Channel verified (No scaling) -0.05 -11.138***
    Party-aligned comment proportion (Exponential) -0.05 -9.771***
    Party-aligned comment proportion (Linear) -0.052 -9.84***
    Party-aligned comment proportion (No scaling) -0.049 -9.735***
    Opposition comment proportion (Exponential) -0.051 -8.608***
    Opposition comment proportion (Linear) -0.051 -8.622***
    Opposition comment proportion (No scaling) -0.051 -8.623***
@@
    Supplementary Table 5: The ideological skew observed in the experiments, as well as the ideological skew computed through counterfactual models which sample videos based on a given attribute. The right-most column denotes independent t-test results between the expected ideological skew and that observed by the bots. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
    Comment Count (Exponential) 0.027 0.626
    Comment Count (Linear) 0.017 0.061
    Comment Count (No scaling) 0.071 3.364**
    Comments per recommendation (Exponential) -0.062 -5.025***
    Comments per recommendation (Linear) -0.067 -5.312***
    Comments per recommendation (No scaling) -0.046 -4.36***
    Combined Author Score (Exponential) -0.078 -5.762***
    Combined Author Score (Linear) -0.078 -6.041***
    Combined Author Score (No scaling) -0.069 -5.37***
    Combined Likes, Shares, Plays Score (Exponential) 0.043 1.418
    Combined Likes, Shares, Plays Score (Linear) 0.038 1.122
    Combined Likes, Shares, Plays Score (No scaling) 0.089 4.159***
    Combined Video Score (Exponential) 0.003 -0.788
    Combined Video Score (Linear) -0.022 -2.227*
    Combined Video Score (No scaling) 0.03 1.053
    Like Count (Exponential) 0.034 0.965
    Like Count (Linear) 0.029 0.652
    Like Count (No scaling) 0.079 3.577***
    Video Length (Exponential) -0.065 -5.162***
    Video Length (Linear) -0.082 -5.796***
    Video Length (No scaling) -0.068 -6.513***
    (Likes + Comments + Shares) / Plays (Exponential) -0.047 -4.074***
    (Likes + Comments + Shares) / Plays (Linear) -0.06 -4.636***
    (Likes + Comments + Shares) / Plays (No scaling) -0.031 -3.711***
    Channel followers (Exponential) -0.014 -2.082*
    Channel followers (Linear) -0.028 -3.031**
    Channel followers (No scaling) 0.016 0.015
    Full Combined score (Exponential) -0.068 -5.339***
    Full Combined score (Linear) -0.073 -5.758***
    Full Combined score (No scaling) -0.058 -5.057***
    Channel cumul. likes (Exponential) -0.019 -2.206*
    Channel cumul. likes (Linear) -0.033 -3.2**
    Channel cumul. likes (No scaling) 0.008 -0.591
    Likes per recommendation (Exponential) -0.046 -3.989***
    Likes per recommendation (Linear) -0.059 -4.602***
    Likes per recommendation (No scaling) -0.034 -3.938***
    Video plays (Exponential) 0.052 1.87
    Video plays (Linear) 0.048 1.575
    Video plays (No scaling) 0.102 4.728***
    Video shares (Exponential) 0.042 1.329
    Video shares (Linear) 0.035 0.984
    Video shares (No scaling) 0.081 3.841***
    Channel video count (Exponential) 0.025 0.511
    Channel video count (Linear) 0.005 -0.703
    Channel video count (No scaling) 0.058 2.9**
    Channel verified (Exponential) -0.074 -5.459***
    Channel verified (Linear) -0.081 -5.941***
    Channel verified (No scaling) -0.066 -5.011***
    Party-aligned comment proportion (Exponential) 0.018 0.119
    Party-aligned comment proportion (Linear) 0.001 -0.874
    Party-aligned comment proportion (No scaling) 0.02 0.245
    Opposition comment proportion (Exponential) -0.016 -1.62
    Opposition comment proportion (Linear) -0.022 -1.963
    Opposition comment proportion (No scaling) -0.016 -1.653
@@
    Supplementary Table 6: The ideological skew observed in the experiments, as well as the ideological skew computed through counterfactual models which sample videos based on a given attribute when only considering positive-partisanship videos (Pro Democrat or Pro Republican). The right-most column denotes independent t-test results between the expected ideological skew and that observed by the bots. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
    Comment Count (Exponential) -0.033 -17.873***
    Comment Count (Linear) -0.046 -17.981***
    Comment Count (No scaling) -0.027 -17.39***
    Comments per recommendation (Exponential) -0.045 -18.194***
    Comments per recommendation (Linear) -0.054 -18.326***
    Comments per recommendation (No scaling) -0.039 -17.763***
    Combined Author Score (Exponential) -0.052 -18.875***
    Combined Author Score (Linear) -0.058 -19.009***
    Combined Author Score (No scaling) -0.048 -19.066***
    Combined Likes, Shares, Plays Score (Exponential) 0.002 -15.192***
    Combined Likes, Shares, Plays Score (Linear) -0.012 -16.356***
    Combined Likes, Shares, Plays Score (No scaling) 0.021 -14.131***
    Combined Video Score (Exponential) -0.02 -17.043***
    Combined Video Score (Linear) -0.034 -17.445***
    Combined Video Score (No scaling) -0.006 -16.478***
    Like Count (Exponential) -0.002 -15.511***
    Like Count (Linear) -0.015 -16.654***
    Like Count (No scaling) 0.016 -14.501***
    Video Length (Exponential) -0.065 -19.538***
    Video Length (Linear) -0.075 -19.655***
    Video Length (No scaling) -0.068 -20.294***
    (Likes + Comments + Shares) / Plays (Exponential) -0.022 -17.163***
    (Likes + Comments + Shares) / Plays (Linear) -0.035 -17.954***
    (Likes + Comments + Shares) / Plays (No scaling) -0.004 -16.426***
    Channel followers (Exponential) -0.033 -18.069***
    Channel followers (Linear) -0.041 -18.175***
    Channel followers (No scaling) -0.021 -17.557***
    Full Combined score (Exponential) -0.045 -18.723***
    Full Combined score (Linear) -0.054 -19.052***
    Full Combined score (No scaling) -0.04 -18.874***
    Channel cumul. likes (Exponential) -0.054 -19.356***
    Channel cumul. likes (Linear) -0.059 -19.191***
    Channel cumul. likes (No scaling) -0.054 -19.995***
    Likes per recommendation (Exponential) -0.022 -17.15***
    Likes per recommendation (Linear) -0.035 -18.002***
    Likes per recommendation (No scaling) -0.002 -16.342***
    Video plays (Exponential) 0.003 -15.049***
    Video plays (Linear) -0.012 -16.294***
    Video plays (No scaling) 0.022 -13.918***
    Video shares (Exponential) 0.001 -15.407***
    Video shares (Linear) -0.011 -16.346***
    Video shares (No scaling) 0.018 -14.478***
    Channel video count (Exponential) -0.054 -19.207***
    Channel video count (Linear) -0.061 -18.856***
    Channel video count (No scaling) -0.051 -19.586***
    Channel verified (Exponential) -0.052 -16.929***
    Channel verified (Linear) -0.057 -17.276***
    Channel verified (No scaling) -0.044 -16.432***
    Party-aligned comment proportion (Exponential) -0.068 -15.53***
    Party-aligned comment proportion (Linear) -0.064 -15.275***
    Party-aligned comment proportion (No scaling) -0.068 -15.431***
    Opposition comment proportion (Exponential) -0.063 -13.608***
    Opposition comment proportion (Linear) -0.06 -13.466***
    Opposition comment proportion (No scaling) -0.064 -13.646***
@@
    Supplementary Table 7: The ideological skew observed in the experiments, as well as the ideological skew computed through counterfactual models which sample videos based on a given attribute when only considering negative-partisanship videos (Anti Democrat or Anti Republican). The right-most column denotes independent t-test results between the expected ideological skew and that observed by the bots. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
    Weight = 0.55 (Exponential) -0.058 -11.238***
    Weight = 0.55 (Linear) -0.065 -11.549***
    Weight = 0.55 (No scaling) -0.049 -10.881***
    Weight = 0.6 (Exponential) -0.058 -11.25***
    Weight = 0.6 (Linear) -0.065 -11.553***
    Weight = 0.6 (No scaling) -0.05 -10.888***
    Weight = 0.65 (Exponential) -0.059 -11.259***
    Weight = 0.65 (Linear) -0.065 -11.537***
    Weight = 0.65 (No scaling) -0.05 -10.9***
    Weight = 0.7 (Exponential) -0.058 -11.244***
    Weight = 0.7 (Linear) -0.065 -11.558***
    Weight = 0.7 (No scaling) -0.049 -10.87***
    Weight = 0.75 (Exponential) -0.058 -11.237***
    Weight = 0.75 (Linear) -0.065 -11.546***
    Weight = 0.75 (No scaling) -0.049 -10.873***
    Weight = 0.8 (Exponential) -0.058 -11.238***
    Weight = 0.8 (Linear) -0.065 -11.557***
    Weight = 0.8 (No scaling) -0.05 -10.908***
    Weight = 0.85 (Exponential) -0.058 -11.248***
    Weight = 0.85 (Linear) -0.065 -11.532***
    Weight = 0.85 (No scaling) -0.049 -10.879***
    Weight = 0.9 (Exponential) -0.058 -11.237***
    Weight = 0.9 (Linear) -0.065 -11.552***
    Weight = 0.9 (No scaling) -0.049 -10.87***
    Weight = 0.95 (Exponential) -0.058 -11.243***
    Weight = 0.95 (Linear) -0.065 -11.548***
    Weight = 0.95 (No scaling) -0.05 -10.905***
    Weight = 1.0 (Exponential) -0.058 -11.256***
    Weight = 1.0 (Linear) -0.065 -11.548***
    Weight = 1.0 (No scaling) -0.05 -10.89***
@@
    Supplementary Table 8: The ideological skew observed in the experiments, as well as the ideological skew computed through counterfactual models which sample videos based on channel verification status with varying weights given to a channels verification status. The right-most column denotes independent t-test results between the expected ideological skew and that observed by the bots. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
    Supplementary Table 9: The 20 most common hashtags for Democrat and Republican conditioning videos.
@@
    Supplementary Table 10: The proportion of Democrat and Republican videos on a given topic, the difference in these proportions, and chi-squared tests comparing these proportions. (* : $p<0.05$, **∗ ∗ : $p<0.01$, ***∗ ∗ ∗ : $p<0.001$)
@@
    Opposition comment proportion 1.2853
    Video Duration 1.0427
    Supplementary Table 11: VIF values for control variables in logistic regression considering all videos viewed during experiment.
@@
    Opposition comment proportion 1.2853
    Video Duration 1.0427
    Supplementary Table 12: VIF values for control variables in logistic regression when considering top account status.
    Variable Coefficient Odds Ratio OR Lower CI OR Upper CI P-value
    Bot conditioning (ref = Rep)
    Democrat 1.0314 2.8051 2.0177 3.8997 p <0.001
    State (ref = NY)
    Georgia -0.0148 0.9853 0.7867 1.234 p = 0.8975
    Texas 0.0524 1.0538 0.839 1.3236 p = 0.6522
    Time
    Week of experiment -0.0477 0.9534 0.9376 0.9695 p <0.001
    Conditioning metrics
    Engagement score of conditioning videos -0.1296 0.8784 0.7506 1.028 p = 0.1062
    Number of conditioning videos -0.0747 0.928 0.8383 1.0273 p = 0.1497
    Recommendation video metrics
    Share count -0.0727 0.9299 0.8309 1.0406 p = 0.2054
    Likes per recommendation -0.2692 0.764 0.6781 0.8608 p <0.001
    Comments per recommendation 0.0591 1.0609 0.9001 1.2505 p = 0.481
    Opposing comment proportion 0.704 2.0219 1.8348 2.2281 p <0.001
    Video length -0.1549 0.8565 0.6763 1.0848 p = 0.1988
    Recommendation channel metrics
    Verified channel -0.5568 0.5731 0.451 0.7282 p <0.001
    Video count 0.1387 1.1487 1.0366 1.273 p = 0.0081
    Follower count 0.1028 1.1082 0.9794 1.254 p = 0.103
    Supplementary Table 13: Logistic regression coefficients and odds-ratios predicting whether a given video watch is mismatched/cross-party aligned.
@@
    Week of experiment -0.04 0.9608 0.9447 0.9771 p <0.001
    Top account x Bot conditioning
    Top account × Democrat bot -1.0949 0.3346 0.2494 0.4489 p <0.001
    Top account × Republican bot 0.1143 1.1211 0.7686 1.6353 p = 0.5527
    Conditioning metrics
    Engagement score of conditioning videos -0.115 0.8914 0.7609 1.0442 p = 0.1545
    Number of conditioning videos -0.072 0.9306 0.8391 1.032 p = 0.1728
    Recommendation video metrics
    Share count -0.052 0.9494 0.861 1.0468 p = 0.297
    Likes per recommendation -0.2392 0.7873 0.6967 0.8896 p <0.001
    Comments per recommendation 0.0488 1.05 0.8887 1.2406 p = 0.5663
    Opposing comment proportion 0.7433 2.1029 1.9043 2.3222 p <0.001
    Video length -0.2003 0.8185 0.6097 1.0988 p = 0.1826